OpenStudy (uri):

A bag contains 6 white and 8 black balls.If 4 balls are drawn.What is the probability that: a)All are black b)2 are white and 2 are black

OpenStudy (uri):

@nincompoop

OpenStudy (bahrom7893):

Sooo you have 14 balls. a) You have to pick all black: (8/14)*(7/13)*(6/12)*(5/11)

OpenStudy (uri):

Got it,and after this? :)

OpenStudy (uri):

@bahrom7893

OpenStudy (bahrom7893):

b) 2 are white and 2 are black: (6/14)*(5/13)*(8/12)*(7/11)

OpenStudy (bahrom7893):

basically probability of picking a ball of a certain color: number of balls of that color / number of balls total

OpenStudy (anonymous):

@bahrom7893 it never said you necessarily draw 2 white and then 2 black!

OpenStudy (uri):

I don't get the b part.

OpenStudy (bahrom7893):

@oldrin.bataku is right. you have to multiply my answer for b by 4!

OpenStudy (bahrom7893):

but wait two of them are the same color, so 4!/2!

OpenStudy (bahrom7893):

I meant to say 2 pairs are the same color

OpenStudy (uri):

Okay now can you solve and show the first part? That is part a. wait we have to write it like this: \[\frac{ 6 }{ 14 }\frac{ 5 }{ 13 }\frac{ 8 }{ 12 }\frac{ 7 }{ 11 }\]

OpenStudy (uri):

Multiply sign in between?

OpenStudy (nincompoop):

how come I didn't get a notification?

OpenStudy (bahrom7893):

yea because basically we have to keep getting black balls: 8 black balls out of 14 total = 8/14 (one black ball is gone) 7 black balls out of 13 total = 7/13, (7 black balls are remaining, and there are 13 in total) etc. follow the pattern

OpenStudy (uri):

I swear to god i don't get it.

OpenStudy (uri):

I miss...@terenzreignz Now..

OpenStudy (uri):

@terenzreignz

OpenStudy (kirbykirby):

You can use the hypergeometric distribution

OpenStudy (kirbykirby):

Let X = # of black balls 14 balls total You choose 4 of them If you want all of them to be black, then plug in x = 4 \[f(x)=\frac{\left(\begin{matrix}8 \\ x\end{matrix}\right) \left(\begin{matrix}6 \\ 4-x\end{matrix}\right)}{\left(\begin{matrix}14 \\ 4\end{matrix}\right)}\]

OpenStudy (kirbykirby):

b) you can just apply the same logic but you consider the case for 2 black balls (put in x=2) then just multiply that with the result for white=2 ... you can re-write f(x) for white balls and compute it for that,

OpenStudy (anonymous):

@kirbykirby is correct; hypergeometric distribution is for where we don't have replacement